[clas_members] [Revised Logentry] RC Daily summary -- Tuesday

daria at jlab.org daria at jlab.org
Tue Sep 13 12:00:02 EDT 2022

Logentry Text:
Dear all, 

*** First of all, a note to all shift-takers: please review the Wiki instructions before every shift, things to monitor / look out for do sometimes change from day to day! ***

The past day's been fairly uneventful (in a good way): taking polarised ND3 data, the monotony punctuated by occasional MVT trips, RICH tantrums, beam disappearances -- and some dirty tricks from MON12.

MVT trips erratically, but seemed particularly nervous in the night (8 trips recorded on the owl shift). 

DAQ had a crash yesterday after lunch, needing Sergey’s intervention.

RICH needed to be “recovered” twice, once in the day, once in the night, but that fixed it.

MON12 has been an evil trickster — when clearing the histograms it sometimes doesn’t reconnect fully: still fills histograms but not all at the same rate, and not in an obvious way. Led us on some wild goose chases.

At Nathan's behest, we have added some more instructions for shift workers regarding RasterMon.

We re-centred the raster a tiny bit: lifted the y-offset by 0.2mm, slight improvement in the raster pattern. https://logbooks.jlab.org/entry/4040611

Lost beam for about 30min this morning, otherwise it’s been OK.

Harut’s analysis shows that after the anneal the NMR agrees again with the polarisation that he extracts from the asymmetry off-line. Thanks for the sustained near-instantaneous cooking and analysing, Mohammad and Harut!

Pol. has been steady (bar beam-down periods) at -30.5% on NMR since 7pm last night. I tweaked the polarising microwave freq. at noon yesterday and again late in the evening — if any improvement, it’s very marginal.

The target doesn’t like much being negative — max polarisation reached (NMR) is -30.5% (it was about 37% for positive pol.) and it took a longer time to get there: we finished annealing in the middle of the day on Sunday, pol. was still only -27% on Monday morning (but it had stalled flat for 6h while beam was off in the night). Reached -30% around 2.30pm on Monday (mid-run 16958). So we have been taking data at peak polarisation for about 22h, according to NMR. 

Tomorrow is a long maintenance day — the beam will go off at 6.30am, they hope to be done by about 6 - 7pm. We will go into restricted access (around 8am).

We suspect that the target doesn’t have much more peak pol. to give today and we need to finish the empty run studies, so swapping to an empty target cell at 5pm (raster with and without LHe).

All the shift workers have been remote for the past day, there we no problems (less company but more snacks for the experts :-).

RC meeting at 2pm today -- might be a short one!

Your still-quite-happy RC

******** Stats since the start of ND3 with ELMO ********

Target irradiation, anneal.

Half-wave plate OUT, target polarisation POSITIVE, good-run triggers:  467 M (16902 - 16908)

Half-wave plate IN, target polarisation POSITIVE, good-run triggers: 498 M (16910 - 16922)

Half-wave plate IN, target polarisation NEGATIVE, good-run triggers: 638 M (16923 - 16832)

Half-wave plate OUT, target polarisation NEGATIVE, good-run triggers: 335 M (16933 - 16944)

--- details in past RC summaries.

Target anneal.

Half-wave plate OUT, target polarisation NEGATIVE:

16945: 78.4 M (raster optimisation, pol. increasing from 12% to 19% -- prob. not for production).
16946: 26.6 M
16947: 67.3 M (DAQ crashed at end).
16948: 17.8 M
16951: 100.2 M (Pinned. NMR about -28% at the end).
16952: 100.1 M
16958: 101.7 M (shifted raster centre up by 0.2 mm around 21.8 M events, not ended run correctly -- but should still be a good one).
16959: 100.0 M (pinned. Peak pol (-30.5%)).

TOTAL triggers from "good" runs (NMR pol > 20%): 514 M.

Half-wave plate IN, target polarisation NEGATIVE:

16960: 77.2 M (DAQ crash at end)
16961: 5.2 M  (pinned. Stopped for RICH prob.)
16963: 100.3 M (pinned)
16964: 100.1 M
16965: running... currently 90 M.

TOTAL triggers from "good" runs so far, all at peak pol (NMR -30.5%): 373 M.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4040992
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/clas_members/attachments/20220913/5ac10027/attachment.html>

More information about the clas_members mailing list